OpenStudy (anonymous):

Solve |2x - 5| = 4. {x | x = -4.5 or x = 4.5} {x | x = 0.5 or x = 4.5} {x | 0.5 < x < 4.5}

OpenStudy (cwrw238):

try plugging in values to see if they fit or you can solve the equation

OpenStudy (cwrw238):

|2x - 5| = 4 means that either 2x - 5 = 4 or 2x - 5 = -4

OpenStudy (anonymous):

is the answer a?

OpenStudy (cwrw238):

no -4.5 is not a root

OpenStudy (cwrw238):

2x - 5 = -4 add 5 to both sides 2x = 1 x = 0.5

OpenStudy (anonymous):

ooh, i see, so it has to be b?

OpenStudy (anonymous):

or am i wrong?

OpenStudy (cwrw238):

yes b

OpenStudy (cwrw238):

the first eqaution gives x = 4.5